# From Objects to Events: Unlocking Complex Visual Understanding in Object Detectors via LLM-guided Symbolic Reasoning

### v3

@@ Abstract
Current object detectors excel at entity localization and classification, yet exhibit inherent limitations in event recognition capabilities. This deficiency arises from their architecture’s emphasis on discrete object identification rather than modeling the compositional reasoning, inter-object correlations, and contextual semantics essential for comprehensive event understanding. To address this challenge, we present a novel framework that expands the capability of standard object detectors beyond mere object recognition to complex event understanding through LLM-guided symbolic reasoning. Our key innovation lies in bridging the semantic gap between object detection and event understanding without requiring expensive task-specific training. The proposed plug-and-play framework interfaces with any open-vocabulary detector while extending their inherent capabilities across architectures. At its core, our approach combines (i) a symbolic regression mechanism exploring relationship patterns among detected entities and (ii) a LLM-guided strategically guiding the search toward meaningful expressions. These discovered symbolic rules transform low-level visual perception into interpretable event understanding, providing a transparent reasoning path from objects to events with strong transferability across domains. We compared our training-free framework against specialized event recognition systems across diverse application domains. Experiments demonstrate that our framework enhances multiple object detector architectures to recognize complex events such as illegal fishing activities (75% AUROC, +8.36% improvement), construction safety violations (+15.77%), and abnormal crowd behaviors (+23.16%). Code is available at here.
@@ 1 Introduction
Object detection has become a cornerstone of computer vision, enabling machines to identify and locate entities within visual scenes with remarkable accuracy [15, 14, 47, 46, 2, 3, 63, 61, 21, 34, 40]. State-of-the-art detectors can now recognize thousands of object categories with remarkable precision, transforming applications across autonomous driving, surveillance, industrial inspection, and content analysis. However, despite these advances, a fundamental limitation persists: while modern detectors excel at answering “what is present” in an image, they struggle with understanding “what is happening” — the relationships, interactions, and events occurring between detected entities.
Consider a coastal surveillance scenario where a detector identifies “persons” and “fishing rods” with high confidence. Despite perfect detection, the system fails to recognize the critical event of “illegal multi-rod fishing” — where a single individual operates multiple fishing rods, violating conservation regulations. Similarly, in construction site monitoring, a standard detector might accurately identify workers, equipment, and safety gear, yet remain incapable of recognizing the crucial safety violation where a worker operates machinery without proper protective equipment. These limitations stem from the architectural focus of object detectors on identifying discrete entities rather than modeling the compositional logic, relational semantics, and contextual dependencies that define meaningful events. Without the ability to understand “what is happening” beyond “what exists,” even the most accurate detection systems fall short in scenarios requiring nuanced interpretation of object relationships and contextual significance — a fundamental barrier to deploying truly intelligent visual systems in complex real-world scenarios.
@@ 2.1 Visual Reasoning and Neuro-symbolic
Our work extends beyond conventional object detection to enable reasoning about complex visual events, placing it within the broader visual reasoning paradigm. While we leverage object detector outputs as our foundation, we transform these into symbolic representations suitable for higher-order reasoning. Visual reasoning research has progressed from simple object recognition to complex scene understanding requiring compositional analysis. Traditional approaches have relied on specialized architectures and extensive labeled datasets for specific reasoning tasks. Visual question answering systems interpret images through natural language questions [20, 4, 22, 10, 38, 5], while scene graph generation approaches identify object relationships to construct structured scene representations [23, 28, 57, 31, 60]. However, these methods often lack interpretability or require task-specific training data. Neuro-symbolic models offer a promising direction by combining neural networks’ perceptual strengths with symbolic reasoning’s interpretability and compositionality [11, 1, 39, 52, 59]. These approaches typically extract symbolic representations from visual scenes using neural networks, then apply symbolic reasoning methods to these representations. Our framework advances this paradigm by implementing a novel neuro-symbolic approach where object detectors serve as the neural perception component while a symbolic reasoning layer guided by LLMs performs higher-level event understanding. Unlike traditional implementations requiring custom integration between components, our approach treats existing object detectors as modular perception units, maintaining interpretability while enabling flexible application across visual domains without task-specific training.
@@ 2.2 LLMs for Visual Tasks
Our approach uniquely positions LLMs as reasoning guides for symbolic search rather than for direct visual perception. This design allows us to leverage LLMs’ rich world knowledge while maintaining a clear separation between perception (via object detectors) and reasoning (via interpretable symbolic operations). Recent advances in LLMs have demonstrated impressive capabilities in visual understanding [45, 37, 17, 54, 61, 29, 65, 51, 6]. Models such as GPT-4V offer high accuracy in complex visual reasoning with low hallucination rates,making them suitable for complex visual analysis and general-purpose visual AI applications [58]. Despite these advances, directly applying LLMs to visual reasoning presents challenges due to modality gaps and reasoning complexity. Our framework addresses these challenges by using LLMs in their native text domain to guide symbolic pattern discovery over detector outputs. This approach maintains complete interpretability throughout the process — a critical advantage over end-to-end black-box models. By separating perception from reasoning, we combine neural models’ perceptual capabilities with symbolic reasoning’s interpretability, enhanced by LLMs’ semantic understanding, without requiring extensive multimodal training.
@@ 3 Method
Figure 2: Illustration of the proposed SymbolicDet. SymbolicDet mainly consists of logic search and symbolic reasoning. The former module constructs and explores the search space by leveraging structured entity features extracted from an open-set object detector. Th latter module harnesses the symbolic reasoning capabilities of Large Language Models (LLMs) along with their inherent commonsense understanding of visual event patterns to guide the search process toward more appropriate and rational pathways.
In practical applications, merely detecting objects often fails to satisfy real-world engineering requirements. Many scenarios demand recognition of complex object relationships or events, which remains a significant challenge in current research. Construction site monitoring requires identifying not just workers and equipment but safety violations; traffic analysis needs to recognize not only vehicles but also dangerous driving patterns; and surveillance systems must detect not merely people but suspicious behaviors. While specialized event recognition systems exist, they typically require extensive training data and lack interpretability. Our framework addresses this challenge by unlocking the latent event understanding capabilities in standard object detectors through LLM-guided symbolic reasoning. Here, we provide a detailed description of our approach, which transforms object detections into interpretable event recognition without additional training. Figure 5 illustrates our framework’s architecture.
@@
where each pair consists of an image $I_{i}$ and a binary label $y_{i}\in\{0,1\}$ indicating whether a target event $\varepsilon$ occurs in the image. More specifically, for each image $I_{i}$, a standard object detector $D$ produces a set of detections $\mathcal{O}=\{o_{1},o_{2},\ldots,o_{m}\}$, where each detection $o_{j}=(c_{j},b_{j},s_{j})$ consists of a category label $c_{j}$, a bounding box $b_{j}$, and a confidence score $s_{j}$. These detections represent “what exists” in the image. Our approach seeks to discover an interpretable symbolic expression $E$ that operates solely on the detector outputs to recognize the event.
@@
$E:\mathcal{O}_{I}\rightarrow\{0,1\}$ (2)
This symbolic expression effectively bridges the gap between low-level object detections and high-level event understanding, transforming “what exists” into “what is happening” through logical reasoning over detected entities and their relationships.These components work together in a synergistic manner. The overall workflow can be represented as:
$E^{*}=\arg\max_{E\in\mathcal{L}}\mathcal{G}_{LLM}(E,S(E,\mathcal{F}(\mathcal{D})))$ (3)
Where $E^{*}$ is the optimal discovered symbolic expression, $\mathcal{L}$ is the space of all possible expressions in our symbolic language, $\mathcal{F}$ is the object detector, $S$ is a scoring function that evaluates how well an expression distinguishes positive and negative examples, and $\mathcal{G}_{LLM}$ is the LLM guidance mechanism that directs the search toward promising expressions. In the following sections, we detail each component of our framework and how they work together to unlock event understanding capabilities in standard object detectors.
@@ 3.1 Symbolic Logic Search
Entity-level Feature Extraction. We first leverage an open-vocabulary object detector to extract comprehensive entity information from each sample. For a given image $x$, we obtain a set of entities:
$E=\{e_{1},e_{2},...,e_{n}\},\quad e_{i}=(c_{i},b_{i},s_{i})$ (4)
where $c_{i}$ represents the category label, $b_{i}=(x,y,w,h)$ denotes the bounding box coordinates, and $s_{i}$ is the detection confidence score. These entities form the basis for our symbolic pattern analysis. To facilitate symbolic reasoning, we transform the raw entity information into structured features:
$\mathbf{X}=\{\phi_{1}(E),\phi_{2}(E),...,\phi_{d}(E)\}$ (5)
@@ 4.1.1 Self-collected Datasets
Helmet-Mac Dataset, which we make publicly available [^1] Dataset will soon be available in our code repository., addresses the critical domain of construction site safety monitoring. This dataset is curated from various construction scenarios and focuses on safety helmet compliance detection. It contains 7,571 training images with detailed annotations of human heads and safety helmets across diverse construction environments. The test set comprises 4,642 images, balanced between 2,276 safety violations (workers without helmets) and 2,366 compliant cases. The dataset captures various challenging scenarios including different lighting conditions, viewing angles, and occlusion cases, making it a valuable benchmark for safety-critical event detection systems.
@@ 4.2 Main Results
Comparison with present Methods. To contextualize our approach within the broader landscape of event detection methods, we evaluate on the UCSD Ped2 benchmark and compare against state-of-the-art approaches. As showed in Table 3, our method achieves an impressive 98.7% accuracy without utilizing semantic-level annotations or task-specific fine-tuning, approaching the state-of-the-art performance (99.7%). This minimal performance gap is particularly interesting considering the vast difference in approach complexity. We hypothesize that this effectiveness stems from two factors: first, the pre-trained detectors already possess rich semantic understanding that generalizes well to visual event pattern; second, our symbolic reasoning framework effectively translates this semantic knowledge into explicit detection rules, potentially capturing patterns that are similar to those learned by supervised methods but in a more interpretable manner.
@@
Table 3: The overall performance on the UCSD ped2 [56] benchmark.This intuitively reflects that the performance of SymbolicDet (APE) without training is very close to the current SOTA.

### v4

@@ Abstract
Current object detectors excel at entity localization and classification, yet exhibit inherent limitations in event recognition capabilities. This deficiency arises from their architecture’s emphasis on discrete object identification rather than modeling the compositional reasoning, inter-object correlations, and contextual semantics essential for comprehensive event understanding. To address this challenge, we present a novel framework that expands the capability of standard object detectors beyond mere object recognition to complex event understanding through LLM-guided symbolic reasoning. Our key innovation lies in bridging the semantic gap between object detection and event understanding without requiring expensive task-specific training. The proposed plug-and-play framework interfaces with any open-vocabulary detector while extending their inherent capabilities across architectures. At its core, our approach combines (i) a symbolic regression mechanism exploring relationship patterns among detected entities and (ii) a LLM-guided strategy guiding the search toward meaningful expressions. These discovered symbolic rules transform low-level visual perception into interpretable event understanding, providing a transparent reasoning path from objects to events with strong transferability across domains. We compared our training-free framework against specialized event recognition systems across diverse application domains. Experiments demonstrate that our framework enhances multiple object detector architectures to recognize complex events such as illegal fishing activities (75% AUROC, +8.36% improvement), construction safety violations (+15.77%), and abnormal crowd behaviors (+23.16%). Code is available at here.
@@ 1 Introduction
Object detection has become a cornerstone of computer vision, enabling machines to identify and locate entities within visual scenes with remarkable accuracy [16, 15, 52, 51, 3, 4, 69, 66, 23, 37, 43]. State-of-the-art detectors can now recognize thousands of object categories with remarkable precision, transforming applications across autonomous driving, surveillance, industrial inspection, and content analysis. However, despite these advances, a fundamental limitation persists: while modern detectors excel at answering “what is present" in an image, they struggle with understanding “what is happening" — the relationships, interactions, and events occurring between detected entities.
Consider a coastal surveillance scenario where a detector identifies “persons" and “fishing rods" with high confidence. Despite perfect detection, the system fails to recognize the critical event of “illegal multi-rod fishing" — where a single individual operates multiple fishing rods, violating conservation regulations. Similarly, in construction site monitoring, a standard detector might accurately identify workers, equipment, and safety gear, yet remain incapable of recognizing the crucial safety violation where a worker operates machinery without proper protective equipment. These limitations stem from the architectural focus of object detectors on identifying discrete entities rather than modeling the compositional logic, relational semantics, and contextual dependencies that define meaningful events. Without the ability to understand “what is happening" beyond “what exists," even the most accurate detection systems fall short in scenarios requiring nuanced interpretation of object relationships and contextual significance — a fundamental barrier to deploying truly intelligent visual systems in complex real-world scenarios.
@@ 2.1 Visual Reasoning and Neuro-symbolic
Our work extends beyond conventional object detection to enable reasoning about complex visual events, placing it within the broader visual reasoning paradigm. While we leverage object detector outputs as our foundation, we transform these into symbolic representations suitable for higher-order reasoning. Visual reasoning research has progressed from simple object recognition to complex scene understanding requiring compositional analysis. Traditional approaches have relied on specialized architectures and extensive labeled datasets for specific reasoning tasks. Visual question answering systems interpret images through natural language questions [22, 5, 24, 11, 41, 6], while scene graph generation approaches identify object relationships to construct structured scene representations [25, 31, 62, 34, 65]. However, these methods often lack interpretability or require task-specific training data. Neuro-symbolic models offer a promising direction by combining neural networks’ perceptual strengths with symbolic reasoning’s interpretability and compositionality [12, 2, 42, 57, 64, 29]. These approaches typically extract symbolic representations from visual scenes using neural networks, then apply symbolic reasoning methods to these representations. Our framework advances this paradigm by implementing a novel neuro-symbolic approach where object detectors serve as the neural perception component while a symbolic reasoning layer guided by LLMs performs higher-level event understanding. Unlike traditional implementations requiring custom integration between components, our approach treats existing object detectors as modular perception units, maintaining interpretability while enabling flexible application across visual domains without task-specific training.
@@ 2.2 LLMs for Visual Tasks
Our approach uniquely positions LLMs as reasoning guides for symbolic search rather than for direct visual perception. This design allows us to leverage LLMs’ rich world knowledge while maintaining a clear separation between perception (via object detectors) and reasoning (via interpretable symbolic operations). Recent advances in LLMs have demonstrated impressive capabilities in visual understanding [50, 40, 19, 59, 66, 32, 71, 56, 7]. DetGPT [49] enables open-vocabulary object detection through reasoning, while Groundhog [68] grounds LLMs to holistic segmentation tasks. Despite these advances, directly applying LLMs to visual reasoning presents challenges due to modality gaps and reasoning complexity. Our framework addresses these challenges by using LLMs in their native text domain to guide symbolic pattern discovery over detector outputs. This approach maintains complete interpretability throughout the process — a critical advantage over end-to-end black-box models. By separating perception from reasoning, we combine neural models’ perceptual capabilities with symbolic reasoning’s interpretability, enhanced by LLMs’ semantic understanding, without requiring extensive multimodal training. Unlike automated prompt optimization that uses search to refine LLM inputs [17, 47, 63], we use the LLM to guide an external symbolic search, rather than treating the LLM as the direct target of optimization.
@@ 3 Method
Figure 2: Illustration of the proposed SymbolicDet. SymbolicDet mainly consists of logic search and symbolic reasoning. The former module constructs and explores the search space by leveraging structured entity features extracted from an open-set object detector. The latter module harnesses the symbolic reasoning capabilities of Large Language Models (LLMs) along with their inherent commonsense understanding of visual event patterns to guide the search process toward more appropriate and rational pathways.
In practical applications, merely detecting objects often fails to satisfy real-world engineering requirements. Many scenarios demand recognition of complex object relationships or events, which remains a significant challenge in current research. Construction site monitoring requires identifying not just workers and equipment but safety violations; traffic analysis needs to recognize not only vehicles but also dangerous driving patterns; and surveillance systems must detect not merely people but suspicious behaviors. While specialized event recognition systems exist, they typically require extensive training data and lack interpretability. Our framework addresses this challenge by unlocking the latent event understanding capabilities in standard object detectors through LLM-guided symbolic reasoning. Here, we provide a detailed description of our approach, which transforms object detections into interpretable event recognition without additional training, See appendix for symbol meanings.. Figure 2 illustrates our framework’s architecture.
@@
where each pair consists of an image $I_{i}$ and a binary label $y_{i}\in\{0,1\}$ indicating whether a target event $\varepsilon$ occurs in the image. More specifically, for each image $I_{i}$, a standard object detector $\mathcal{T}$ produces a set of detections $D_{i}=\{d_{1},d_{2},\ldots,d_{m}\}$, where each detection $d_{j}=(c_{j},b_{j},s_{j})$ consists of a category label $c_{j}$, a bounding box $b_{j}$, and a confidence score $s_{j}$. These detections represent “what exists" in the image. Our approach seeks to discover an interpretable symbolic expression $f$ that operates solely on the detector outputs to recognize the event.
@@
$f:\mathcal{O}_{I}\rightarrow\{0,1\}$ (2)
This symbolic expression effectively bridges the gap between low-level object detections and high-level event understanding, transforming “what exists" into “what is happening" through logical reasoning over detected entities and their relationships. These components operate synergistically, constituting an integrated workflow:
$f^{*}=\arg\max_{f\in\mathcal{F}}\mathcal{G}_{LLM}(f,S(f,\mathcal{T}(\mathcal{D})))$ (3)
Where $f^{*}$ is the optimal discovered symbolic expression, $\mathcal{F}$ is the space of all possible expressions in our symbolic language, $\mathcal{T}$ is the object detector, $S$ is a scoring function that evaluates how well an expression distinguishes positive and negative examples, and $\mathcal{G}_{LLM}$ is the LLM guidance mechanism that directs the search toward promising expressions. In the following sections, we detail each component of our framework and how they work together to unlock event understanding capabilities in standard object detectors.
@@ 3.1 Symbolic Logic Search
Entity-level Feature Extraction. We first leverage an open-vocabulary object detector to extract comprehensive entity information from each sample. For a given image $I_{i}$, we obtain a set of entities $D_{i}$. These entities form the basis for our symbolic pattern analysis. To facilitate symbolic reasoning, we transform the raw entity information into structured features:
$\mathbf{X}=\{\phi_{1}(D_{1}),\phi_{2}(D_{2}),...,\phi_{d}(D_{n})\}$ (4)
@@ 4.1.1 Self-collected Datasets
Helmet-Mac Dataset, which we make publicly available***Dataset will soon be available in our code repository., addresses the critical domain of construction site safety monitoring. This dataset is curated from various construction scenarios and focuses on safety helmet compliance detection. It contains 7,571 training images with detailed annotations of human heads and safety helmets across diverse construction environments. The test set comprises 4,642 images, balanced between 2,276 safety violations (workers without helmets) and 2,366 compliant cases. The dataset captures various challenging scenarios including different lighting conditions, viewing angles, and occlusion cases, making it a valuable benchmark for safety-critical event detection systems.
@@ 4.2 Main Results
Comparison with present Methods. To contextualize our approach within the broader landscape of event detection methods, we evaluate on the UCSD Ped2 and USED[1] benchmark and compare against state-of-the-art approaches. As showed in Table 3, our method achieves an impressive 98.7% accuracy without utilizing training is very close to the current SOTA in UCSD Ped2 dataset and outperforms the baseline in USED dataset. This minimal performance gap is particularly interesting considering the vast difference in approach complexity. We hypothesize that this effectiveness stems from two factors: first, the pre-trained detectors already possess rich semantic understanding that generalizes well to visual event pattern; second, our symbolic reasoning framework effectively translates this semantic knowledge into explicit detection rules, potentially capturing patterns that are similar to those learned by supervised methods but in a more interpretable manner.
@@
Table 3: The overall performance on the UCSD ped2 [61] and USED[1] benchmark.
@@ 5 Conclusion
ACKNOWLEDGMENTS
This work was supported by the National Science Fund for Distinguished Young Scholars (No.62025603), the National Natural Science Foundation of China (No. U21B2037, No. U22B2051, No. U23A20383, No. U21A20472, No. 62176222, No. 62176223, No. 62176226, No. 62072386, No. 62072387, No. 62072389, No. 62002305, No. 62272401 and No. 62402015), the Natural Science Foundation of Fujian Province of China (No. 2021J06003, No.2022J06001), the Postdocotoral Fellowship Program of CPSF (GZB20230024) and the China Postdoctoral Science Foundation (2024M750100).
@@ 6 Ablation Study
Ours NSCL NS-VQA LogicHOI
Multi-rods 0.75 0.65 0.61 0.67
Helmet 0.83 0.68 0.65 0.69
Table 4: Comparison with other nuero-symbolic methods. (Acc)
@@
wo llm 7B 14B 32B 72B
Run time(s/500it) 80.66 265.5 281.01 268 267
Cost time(s)1 69.66 45.92 44.91 30.35 26.07
Memory(MB) 293.65 218.56 218.34 218.68 218.82
Table 5: The computational overhead of symbolic search process.
  • 1
    It refers to the time needed to achieve the same performance.
@@
7 Extra experiments
We conducted additional experiments by transferring several representative neuro-symbolic methods to be evaluated on our benchmark dataset. Detailed results can be found in Table 4. In addition, we conducted experiments to compare the computational overhead. We analyze the trade-off between computational cost and performance when using LLMs with varying parameter counts versus not using an LLM at all.Details can be found in Table 5. We also analyze the computational costs of each step of our work.The main computational steps of this work are: Detector inference, Symbolic search process and LLM-guided reasoning. It is worth noting that we use a third-party LLM service provider for the third step of computation, so the actual computational cost may vary significantly depending on the service provider. Detailed computational costs for each part are shown in Table 6.
Detection Search LLm Infer
Run time(s/iter) 0.02-1.5 0.16 8-14
GPU Memory(GB) 0.54-18 - -
Table 6: The computational overhead of each part.
In addition, we have conducted additional experiments on three subsets of the USED dataset: SPORT, CONCERT, and PROTEST. The results from these new experiments will be presented in Table 7.
SPORT CONCERT PROTEST
Ours 0.93 0.99 0.92
USED 0.66 0.75 0.67
Table 7: The computational overhead of each part. (Acc)
@@ 7 Extra experiments
Symbol Definition
$\mathcal{D}$ A visual datasdet
$I_{i}$ An image i in visual dataset
$y_{i}$ A binary label corresponds to image i
${D}$ A set of objects which are detected by object detector
$d_{j}$ A detected object
$c_{j}$ Category label belongs to $d_{j}$
$b_{j}$ Bounding box belongs to $d_{j}$
$s_{j}$ Confidence score belongs to $d_{j}$
$\epsilon$ Target event
$f$ Symbolic expression
$\mathcal{O}_{I}$ A binary symbol to judge whether the target event is present within an image I
$f^{*}$ Optimal discovered symbolic expression
$\mathcal{F}$ The space of all possible expressions in our symbolic language
$\mathcal{T}$ Object detector
$S$ A scoring function that evaluates how well an expression distinguishes positive and negative examples
$\mathcal{G}_{LLM}$ The LLM guidance mechanism that directs the search toward promising expressions
$\phi_{i}(\cdot)$ Feature extraction functions that capture entity counts, spatial relationships, and attribute distributions
$\Omega(f)$ A complexity penalty that promotes simpler expressions
$P_{\text{init}}$ Initial prompt for the LLM guidence
$P_{\text{cot}}$ Prompt for chain-of-thought
$P_{\text{feed}}$ Prompt after contextual feedback integration
Table 8: Symbol Definition
8 Notation and Results
This section provides a comprehensive list of mathematical symbols and notations used throughout this paper, followed by visual demonstration of our proposed method’s performance. Table 8 summarizes the key mathematical symbols and their definitions used in this work. Figure 5 illustrates the effectiveness of our proposed approach through visual comparison and performance metrics.